# Quantum Encrypted Control of Networked Systems

### v2

@@ Abstract
keywords:
Networked systems, encrypted control, quantum entanglements
thanks: A preliminary work has been submitted to the 23rd IFAC World Congress, August 23-28, 2026, BEXCO, Busan, Republic of Korea. This work was supported by the Australian Research Council under Grant DP190103615, Grant LP210200473, and Grant DP230101014, and by the Faculty of Engineering’s Breakthrough Project at The University of Sydney.
@@ 3.2 Quantum Encrypted Control Scheme
$f_{\rm Enc}:\mathbb{R}^{n}\times\{0,1\}_{w_{\rm q}}\rightarrow\mathbb{C}_{x},$
$f_{\rm Con}:\mathbb{R}^{1\times n}\times\mathbb{C}_{x}\rightarrow\mathbb{C}_{u},$
$f_{\rm Dec}:\mathbb{C}_{u}\times\{0,1\}_{w_{\rm q}}\rightarrow\mathbb{R}$
@@ 4 Stability under Intrinsic Quantum Error
Assumption 1 (Quantum Key Errors).
@@ 5.1 QEC with Quantization
$\mathbb{P}\left(\sum_{j=0}^{w-1}2^{-j}a_{j}=\frac{\lfloor 2^{w-1}y\rfloor}{2^{w-1}}\right)$ (10)
$=$ $1-(2^{w-1}y-\lfloor 2^{w-1}y\rfloor),$
$\mathbb{P}\left(\sum_{j=0}^{w-1}2^{-j}a_{j}=\frac{\lfloor 2^{w-1}y\rfloor}{2^{w-1}}+\frac{1}{2^{w-1}}\right)$
$=$ $2^{w-1}y-\lfloor 2^{w-1}y\rfloor;$
@@
Exponential-Logarithm Realization of $\mathsf{QEC}$ with Quantization The parameters $\alpha,\overline{x}_{i},\delta_{i}>0$ for i=1,2,,ni=1,2,...,n are globally shared. From $q(t)$, we define $\beta_{i}(t)$= $-(2^{w_{b}-1}+1){b_{i,w_{b}-1}}(t)+\sum_{j=0}^{w_{b}-2}2^{j}b_{i,j}(t)$ (13) $+1+(-1)^{b_{i,w_{b-1}}(t)}\frac{\overline{x}_{i}}{\ln(1+\alpha)},$ for i=1,2,,ni=1,2,...,n. (i) Encryption function: The function $f_{\rm Enc}$ under which $\tilde{\mathbf{x}}(t)=f_{\rm Enc}(\mathbf{x}(t),q(t))$ is defined by $v_{i}(t)$ $=\exp({[\mathbf{x}(t)]_{i}/\beta_{i}(t)}),$ (14) $[\tilde{\mathbf{x}}]_{i}(t)$ $=\overline{c_{i,w-1}c_{i,w-2}\ldots c_{i,0}}(t)=Q_{w}(v_{i}(t)),$ for i=1,2,,ni=1,2,...,n. (ii) Control law function: The function $f_{\rm Con}$ under which $\tilde{\mathbf{u}}(t)=f_{\rm Con}(\mathbf{K},\tilde{\mathbf{x}}(t))$ is defined by $z_{i}$ $=\left(g^{-1}\left(\sum_{j=0}^{w-1}2^{-j}c_{i,j}\right)\right)^{[\mathbf{K}]_{i}},$ (15) $[\tilde{\mathbf{u}}]_{i}$ $=\overline{d_{i,w-1}d_{i,w-2}\ldots d_{i,0}}=Q_{w}\left(z_{i}^{1/\delta_{i}}\right),$ for i=1,2,,ni=1,2,...,n. (iii) Decryption function: The function $f_{\rm Dec}$ under which $\mathbf{u}(t)=f_{\rm Dec}(\tilde{\mathbf{u}}(t),q(t))$ is defined by $\mathbf{u}$ $=\sum_{i=1}^{n}\delta_{i}{\beta}_{i}\ln\left(g^{-1}\left(\sum_{j=0}^{w-1}2^{-j}d_{i,j}\right)\right).$ (16)
@@ 6.1 Differential Privacy of Quantum Keys
$\beta_{i}(t)\in\mathcal{A}_{\beta}$= $\left[-2^{w_{b}-1}-\frac{\overline{x}_{i}}{\ln(1+\alpha)},-\frac{\overline{x}_{i}}{\ln(1+\alpha)}\right)\cup$
$\left(\frac{\overline{x}_{i}}{\ln(1+\alpha)},2^{w_{b}-1}+\frac{\overline{x}_{i}}{\ln(1+\alpha)}\right]$
@@ Appendix A Proof for Theorem 2
$\mathsf{D}(\lambda_{i}(t))$ $=\mathsf{E}(\mathsf{D}(\lambda_{i}(t)|\beta_{i}(t)))+\mathsf{D}(\mathsf{E}(\lambda_{i}(t)|\beta_{i}(t)))$ (24)
$=\mathsf{E}\left(\frac{4^{w_{b}}+3\cdot 2^{w_{b}}+2}{3}\frac{p-p^{2}}{\beta_{i}^{2}(t)}\right)+0=:h(p)$
@@
$\mathbf{x}(t+1)\mathbf{x}^{\top}(t+1)=$ $\mathbf{M}_{0}\mathbf{x}(t)\mathbf{x}^{\top}(t)\mathbf{M}_{0}+\mathbf{M}_{0}\mathbf{x}(t)\mathbf{w}^{\top}(t)\mathbf{B}^{\top}$
$+\mathbf{B}\mathbf{w}(t)\mathbf{x}^{\top}(t)\mathbf{M}_{0}+\mathbf{B}\mathbf{w}(t)\mathbf{w}^{\top}(t)\mathbf{B}^{\top},$
@@
$=$ $\mathbf{M}_{0}\mathbf{x}(t)\mathbf{x}^{\top}(t)\mathbf{M}_{0}+\mathsf{E}(\mathbf{w}(t)|\mathbf{x}(t))\mathbf{M}_{0}\mathbf{x}(t)\mathbf{B}^{\top}$
$+\mathsf{E}(\mathbf{w}(t)|\mathbf{x}(t))\mathbf{B}\mathbf{x}^{\top}(t)\mathbf{M}_{0}+\mathsf{E}(\mathbf{w}^{2}(t)|\mathbf{x}(t))\mathbf{B}\mathbf{B}^{\top}$
$=$ $\mathbf{M}_{0}\mathbf{x}(t)\mathbf{x}^{\top}(t)\mathbf{M}_{0}+2p\mathbf{M}_{0}\mathbf{x}(t)\mathbf{x}^{\top}(t)\mathbf{M}_{1}^{\top}$
$+2p\mathbf{M}_{1}\mathbf{x}(t)\mathbf{x}^{\top}(t)\mathbf{M}_{0}+\overline{h}(p)\mathbf{M}_{1}\mathbf{x}(t)\mathbf{x}^{\top}(t)\mathbf{M}_{1}^{\top},$
@@
$\mathbf{V}(t+1)=$ $\big(\mathbf{M}_{0}\otimes\mathbf{M}_{0}+2p(\mathbf{M}_{1}\otimes\mathbf{M}_{0}+\mathbf{M}_{0}\otimes\mathbf{M}_{1})$
$+\overline{h}(p)\mathbf{M}_{1}\otimes\mathbf{M}_{1}\big)\mathbf{V}(t).$
@@
$\mathbf{M}(p)$= $\mathbf{M}_{0}\otimes\mathbf{M}_{0}+2p\,(\mathbf{M}_{1}\otimes\mathbf{M}_{0}+\mathbf{M}_{0}\otimes\mathbf{M}_{1})$
$+\overline{h}(p)\,\mathbf{M}_{1}\otimes\mathbf{M}_{1},$
@@ Appendix B Proof for Theorem 3
$\leq$ $2\mathsf{E}\left(\sum_{i=1}^{n}\beta_{i}\ln(z_{i})-\sum_{i=1}^{n}[\mathbf{K}]_{i}[\mathbf{x}]_{i}\right)^{2}$
$+2\mathsf{E}\left(\sum_{i=1}^{n}\delta_{i}\beta_{i}\ln\!\left(\frac{g^{-1}\left(\sum_{j=0}^{w-1}2^{-j}d_{i,j}\right)}{z^{1/\delta}}\right)\right)^{2}.$
@@
$\mathsf{E}\left(g^{-1}\left(\sum_{j=0}^{w-1}2^{-j}a_{j}\right)-v\right)^{2}$ $\leq\mathsf{E}\left(\sum_{j=0}^{w-1}2^{-j}a_{j}-g(v)\right)^{2}$ (30)
$\leq\frac{1}{2^{2w}},$
@@
$z_{i}$ $=\left(v_{i}+e_{i}\right)^{[\mathbf{K}]_{i}}$
$=v_{i}^{[\mathbf{K}]_{i}}\left(1+\frac{e_{i}}{v_{i}}\right)^{[\mathbf{K}]_{i}}$
$=z^{r}_{i}\left(1+\frac{e_{i}}{v_{i}}\right)^{[\mathbf{K}]_{i}}$
@@
$=$ $\sum_{i=1}^{n}\mathsf{E}\left(\beta_{i}\ln\left(1+\frac{z_{i}}{z_{i}^{r}}-1\right)\right)^{2}$
$+2\sum_{i<j}\mathsf{E}\left(\beta_{i}\ln\left(1+\frac{z_{i}}{z_{i}^{r}}-1\right)\right)\mathsf{E}\left(\beta_{j}\ln\left(1+\frac{z_{j}}{z_{j}^{r}}-1\right)\right)$
@@
$\mathsf{E}\left(\sum_{i=1}^{n}\delta_{i}\beta_{i}\ln\!\left(1+\frac{\sum_{j=0}^{w-1}2^{-j}d_{i,j}}{z^{1/\delta_{i}}}-1\right)\right)^{2}$ (36)
$\leq$ $\sum_{i=1}^{n}\frac{(\delta_{i}\beta_{i})^{2}}{2^{2w}}+o\!\left(\frac{1}{2^{2w}}\right),$
@@ Appendix C Proof for Theorem 4
$|v_{i}(t)-v_{i}^{\prime}(t)|$ (38)
$\leq$ $\zeta\sup_{|\beta|\geq\frac{\overline{x}_{i}}{\ln(1+\alpha)}}\frac{{\rm d}(\exp\left(\frac{[\mathbf{x}(t)]_{i}}{\beta}\right))}{{\rm d}\beta}$
$\leq$ $\zeta\frac{(1+\alpha)\ln^{2}(1+\alpha)}{\overline{x}_{i}},$
@@
$|\mathbb{P}(Q_{w}(v_{i}(t))\in\mathbb{S})-\mathbb{P}(Q_{w}(v_{i}^{\prime}(t))\in\mathbb{S})|$ (39)
$\leq$ $\zeta\frac{(1+\alpha)\ln^{2}(1+\alpha)}{\overline{x}_{i}}{2^{w-1}},\quad\forall\ \mathbb{S}\in\mathcal{F}_{0}.$
@@
$\sup_{\mathcal{\mathbb{O}\in\mathcal{F}}}|\mathbb{P}(O_{T}(\mathbf{E}_{T})\in\mathbb{O})-\mathbb{P}(O_{T}(\mathbf{E}_{T}^{\prime}))\in\mathbb{O})|$
$\leq$ $\sup_{\mathcal{\mathbb{S}\in\mathcal{F}}_{0}}|\mathbb{P}([\tilde{\mathbf{x}}_{i^{\ast}}(t^{\ast})]\in\mathbb{S})-\mathbb{P}([\tilde{\mathbf{x}}_{i^{\ast}}^{\prime}(t^{\ast})]\in\mathbb{S})|$
$\leq$ $\zeta\frac{(1+\alpha)\ln^{2}(1+\alpha)}{\overline{x}_{i^{\ast}}}{2^{w-1}},$

### v3

@@ Abstract
Keywords: Networked systems, encrypted control, quantum entanglements
@@ 3.2 Quantum Encrypted Control Scheme
$f_{\rm Enc}:\mathbb{R}^{n}\times\{0,1\}_{w_{\rm q}}\rightarrow\mathbb{C}_{x},\quad f_{\rm Con}:\mathbb{R}^{1\times n}\times\mathbb{C}_{x}\rightarrow\mathbb{C}_{u},\quad f_{\rm Dec}:\mathbb{C}_{u}\times\{0,1\}_{w_{\rm q}}\rightarrow\mathbb{R}$
@@ 4 Stability under Intrinsic Quantum Error
Assumption 1 (Quantum Key Errors)
@@ 5.1 QEC with Quantization
$\mathbb{P}\left(\sum_{j=0}^{w-1}2^{-j}a_{j}=\frac{\lfloor 2^{w-1}y\rfloor}{2^{w-1}}\right)=1-(2^{w-1}y-\lfloor 2^{w-1}y\rfloor),$ (10)
$\mathbb{P}\left(\sum_{j=0}^{w-1}2^{-j}a_{j}=\frac{\lfloor 2^{w-1}y\rfloor}{2^{w-1}}+\frac{1}{2^{w-1}}\right)=2^{w-1}y-\lfloor 2^{w-1}y\rfloor;$
@@
Exponential-Logarithm Realization of $\mathsf{QEC}$ with Quantization The parameters $\alpha,\overline{x}_{i},\delta_{i}>0$ for i=1,2,,ni=1,2,...,n are globally shared. From $q(t)$, we define $\beta_{i}(t)$= $-(2^{w_{b}-1}+1){b_{i,w_{b}-1}}(t)+\sum_{j=0}^{w_{b}-2}2^{j}b_{i,j}(t)+1+(-1)^{b_{i,w_{b-1}}(t)}\frac{\overline{x}_{i}}{\ln(1+\alpha)},$ (13) for i=1,2,,ni=1,2,...,n. (i) Encryption function: The function $f_{\rm Enc}$ under which $\tilde{\mathbf{x}}(t)=f_{\rm Enc}(\mathbf{x}(t),q(t))$ is defined by $v_{i}(t)$ $=\exp({[\mathbf{x}(t)]_{i}/\beta_{i}(t)}),$ (14) $[\tilde{\mathbf{x}}]_{i}(t)$ $=\overline{c_{i,w-1}c_{i,w-2}\ldots c_{i,0}}(t)=Q_{w}(v_{i}(t)),$ for i=1,2,,ni=1,2,...,n. (ii) Control law function: The function $f_{\rm Con}$ under which $\tilde{\mathbf{u}}(t)=f_{\rm Con}(\mathbf{K},\tilde{\mathbf{x}}(t))$ is defined by $z_{i}$ $=\left(g^{-1}\left(\sum_{j=0}^{w-1}2^{-j}c_{i,j}\right)\right)^{[\mathbf{K}]_{i}},$ (15) $[\tilde{\mathbf{u}}]_{i}$ $=\overline{d_{i,w-1}d_{i,w-2}\ldots d_{i,0}}=Q_{w}\left(z_{i}^{1/\delta_{i}}\right),$ for i=1,2,,ni=1,2,...,n. (iii) Decryption function: The function $f_{\rm Dec}$ under which $\mathbf{u}(t)=f_{\rm Dec}(\tilde{\mathbf{u}}(t),q(t))$ is defined by $\mathbf{u}$ $=\sum_{i=1}^{n}\delta_{i}{\beta}_{i}\ln\left(g^{-1}\left(\sum_{j=0}^{w-1}2^{-j}d_{i,j}\right)\right).$ (16)
@@ 6.1 Differential Privacy of Quantum Keys
$\beta_{i}(t)\in\mathcal{A}_{\beta}$= $\left[-2^{w_{b}-1}-\frac{\overline{x}_{i}}{\ln(1+\alpha)},-\frac{\overline{x}_{i}}{\ln(1+\alpha)}\right)\cup\left(\frac{\overline{x}_{i}}{\ln(1+\alpha)},2^{w_{b}-1}+\frac{\overline{x}_{i}}{\ln(1+\alpha)}\right]$
@@ Appendix A Proof for Theorem 2
$\mathsf{D}(\lambda_{i}(t))$ $=\mathsf{E}(\mathsf{D}(\lambda_{i}(t)|\beta_{i}(t)))+\mathsf{D}(\mathsf{E}(\lambda_{i}(t)|\beta_{i}(t)))=\mathsf{E}\left(\frac{4^{w_{b}}+3\cdot 2^{w_{b}}+2}{3}\frac{p-p^{2}}{\beta_{i}^{2}(t)}\right)+0=:h(p)$ (24)
@@
$\mathbf{x}(t+1)\mathbf{x}^{\top}(t+1)=$ $\mathbf{M}_{0}\mathbf{x}(t)\mathbf{x}^{\top}(t)\mathbf{M}_{0}+\mathbf{M}_{0}\mathbf{x}(t)\mathbf{w}^{\top}(t)\mathbf{B}^{\top}+\mathbf{B}\mathbf{w}(t)\mathbf{x}^{\top}(t)\mathbf{M}_{0}+\mathbf{B}\mathbf{w}(t)\mathbf{w}^{\top}(t)\mathbf{B}^{\top},$
@@
$=$ $\mathbf{M}_{0}\mathbf{x}(t)\mathbf{x}^{\top}(t)\mathbf{M}_{0}+\mathsf{E}(\mathbf{w}(t)|\mathbf{x}(t))\mathbf{M}_{0}\mathbf{x}(t)\mathbf{B}^{\top}+\mathsf{E}(\mathbf{w}(t)|\mathbf{x}(t))\mathbf{B}\mathbf{x}^{\top}(t)\mathbf{M}_{0}+\mathsf{E}(\mathbf{w}^{2}(t)|\mathbf{x}(t))\mathbf{B}\mathbf{B}^{\top}$
$=$ $\mathbf{M}_{0}\mathbf{x}(t)\mathbf{x}^{\top}(t)\mathbf{M}_{0}+2p\mathbf{M}_{0}\mathbf{x}(t)\mathbf{x}^{\top}(t)\mathbf{M}_{1}^{\top}+2p\mathbf{M}_{1}\mathbf{x}(t)\mathbf{x}^{\top}(t)\mathbf{M}_{0}+\overline{h}(p)\mathbf{M}_{1}\mathbf{x}(t)\mathbf{x}^{\top}(t)\mathbf{M}_{1}^{\top},$
@@
$\mathbf{V}(t+1)=$ $\big(\mathbf{M}_{0}\otimes\mathbf{M}_{0}+2p(\mathbf{M}_{1}\otimes\mathbf{M}_{0}+\mathbf{M}_{0}\otimes\mathbf{M}_{1})+\overline{h}(p)\mathbf{M}_{1}\otimes\mathbf{M}_{1}\big)\mathbf{V}(t).$
@@
$\mathbf{M}(p)$= $\mathbf{M}_{0}\otimes\mathbf{M}_{0}+2p\,(\mathbf{M}_{1}\otimes\mathbf{M}_{0}+\mathbf{M}_{0}\otimes\mathbf{M}_{1})+\overline{h}(p)\,\mathbf{M}_{1}\otimes\mathbf{M}_{1},$
@@ Appendix B Proof for Theorem 3
$\leq$ $2\mathsf{E}\left(\sum_{i=1}^{n}\beta_{i}\ln(z_{i})-\sum_{i=1}^{n}[\mathbf{K}]_{i}[\mathbf{x}]_{i}\right)^{2}+2\mathsf{E}\left(\sum_{i=1}^{n}\delta_{i}\beta_{i}\ln\!\left(\frac{g^{-1}\left(\sum_{j=0}^{w-1}2^{-j}d_{i,j}\right)}{z^{1/\delta}}\right)\right)^{2}.$
@@
$\mathsf{E}\left(g^{-1}\left(\sum_{j=0}^{w-1}2^{-j}a_{j}\right)-v\right)^{2}$ $\leq\mathsf{E}\left(\sum_{j=0}^{w-1}2^{-j}a_{j}-g(v)\right)^{2}\leq\frac{1}{2^{2w}},$ (30)
@@
$z_{i}$ $=\left(v_{i}+e_{i}\right)^{[\mathbf{K}]_{i}}=v_{i}^{[\mathbf{K}]_{i}}\left(1+\frac{e_{i}}{v_{i}}\right)^{[\mathbf{K}]_{i}}=z^{r}_{i}\left(1+\frac{e_{i}}{v_{i}}\right)^{[\mathbf{K}]_{i}}$
@@
$=$ $\sum_{i=1}^{n}\mathsf{E}\left(\beta_{i}\ln\left(1+\frac{z_{i}}{z_{i}^{r}}-1\right)\right)^{2}+2\sum_{i<j}\mathsf{E}\left(\beta_{i}\ln\left(1+\frac{z_{i}}{z_{i}^{r}}-1\right)\right)\mathsf{E}\left(\beta_{j}\ln\left(1+\frac{z_{j}}{z_{j}^{r}}-1\right)\right)$
@@
$\mathsf{E}\left(\sum_{i=1}^{n}\delta_{i}\beta_{i}\ln\!\left(1+\frac{\sum_{j=0}^{w-1}2^{-j}d_{i,j}}{z^{1/\delta_{i}}}-1\right)\right)^{2}\leq\sum_{i=1}^{n}\frac{(\delta_{i}\beta_{i})^{2}}{2^{2w}}+o\!\left(\frac{1}{2^{2w}}\right),$ (36)
@@ Appendix C Proof for Theorem 4
$|v_{i}(t)-v_{i}^{\prime}(t)|\leq\zeta\sup_{|\beta|\geq\frac{\overline{x}_{i}}{\ln(1+\alpha)}}\frac{{\rm d}(\exp\left(\frac{[\mathbf{x}(t)]_{i}}{\beta}\right))}{{\rm d}\beta}\leq\zeta\frac{(1+\alpha)\ln^{2}(1+\alpha)}{\overline{x}_{i}},$ (38)
@@
$|\mathbb{P}(Q_{w}(v_{i}(t))\in\mathbb{S})-\mathbb{P}(Q_{w}(v_{i}^{\prime}(t))\in\mathbb{S})|\leq\zeta\frac{(1+\alpha)\ln^{2}(1+\alpha)}{\overline{x}_{i}}{2^{w-1}},\quad\forall\ \mathbb{S}\in\mathcal{F}_{0}.$ (39)
@@
$\sup_{\mathcal{\mathbb{O}\in\mathcal{F}}}|\mathbb{P}(O_{T}(\mathbf{E}_{T})\in\mathbb{O})-\mathbb{P}(O_{T}(\mathbf{E}_{T}^{\prime}))\in\mathbb{O})|\leq\sup_{\mathcal{\mathbb{S}\in\mathcal{F}}_{0}}|\mathbb{P}([\tilde{\mathbf{x}}_{i^{\ast}}(t^{\ast})]\in\mathbb{S})-\mathbb{P}([\tilde{\mathbf{x}}_{i^{\ast}}^{\prime}(t^{\ast})]\in\mathbb{S})|\leq\zeta\frac{(1+\alpha)\ln^{2}(1+\alpha)}{\overline{x}_{i^{\ast}}}{2^{w-1}},$